### v2

# Kang M, Chen D, Pu Z, Gao J, Yu W. Spatio-Temporal Graphical Counterfactuals: An Overview

Abstract
Counterfactual thinking is a critical yet challenging topic for artificial intelligence to learn knowledge from data and ultimately improve their performances for new scenarios. Many research works, including Potential Outcome Model and Structural Causal Model, have been proposed to realize it. However, their modelings, theoretical foundations and application approaches are usually different. Moreover, there is a lack of graphical approach to infer spatio-temporal counterfactuals, that considers spatial and temporal interactions between multiple units. Thus, in this work, our aim is to investigate a survey to compare and discuss different counterfactual models, theories and approaches, and further build a unified graphical causal framework to infer the spatio-temporal counterfactuals.
keywords:
Counterfactual inference, spatio-temporal graphical model, Potential Outcome Model, Structural Causal Model, causality
1 Introduction
How to enable an intelligent machine to think and answer a spatio-temporal counterfactual question? E.g., sometimes we want to know, if a different investment strategy had been implemented, would we obtain higher returns? Or, in computer networks, what changes would occur in network load if a node’s configuration had never been changed? Or, would someone still purchase the product even if they had never been shown the advertisement? These questions generally have the spatio-temporal counterfactual pattern, that is, “Observed $\dots$, if had done $\dots$, how would $\dots$ go?”. Thus, according to the pattern, firstly, time is necessary to consider, because the questions are queried now, but the counterfactual actions are supposed to be done in the past. Secondly, counterfactual outcomes are unobservable. The outcomes of some actions are observed if the actions have been done to the real-world system, thus these outcomes are factual. But the counterfactual outcomes are not observed, and they are imaged through supposing different actions had been done. Thirdly, the real-world system can be observed as multivariate time series, and there are widespread multivariate interactions within the time series [1, 2].
To enable an intelligent machine to imitate counterfactual thinking like human, thus, Potential Outcome Model (POM) [3] and Structural Causal Model (SCM) [4, 5, 6, 7] are proposed as two frameworks with different foundations. POM uses stable unit treatment value assumption (SUTVA), consistency assumption, positivity assumption and ignorability assumption as foundations [8, 9, 10, 11]. If these assumptions are satisfied, one can use matching or imputation approaches to infer counterfactuals without confounders. But in contrast, SCM builds a causal ladder, i.e., abduction, action and prediction, to infer counterfactuals through graphical language d-separation [4, 5, 7]. With the graphical language d-separation, then causal Markov assumption, faithfulness assumption and causal sufficiency assumption are proposed to guarantee the correctness of counterfactual inference on SCM [12, 13, 14, 15, 16].
However, the different foundations between POM and SCM lead to differences in counterfactual inference. Such differences include: (i) POM assumes the positivity, but SCM does not. That means the counterfactuals in POM are factual actually and can be found in observations certainly. But in contrast, the counterfactuals in SCM can be out of observations and unfalsifiable, that is, one cannot validate the correctness of the inferred counterfactuals. (ii) POM assumes the consistency, but it is rejected for SCM [7]. Actually, SCM also needs the consistency assumption implicitly to guarantee that the counterfactual outcome is one of the unobserved potential outcomes. (iii) SCM uses a type of graphical language to deconfound with certain rules, but POM mainly relies on human-guided deconfounding, that makes it difficult to satisfy the ignorability. But in contrast, SCM can deconfound through back-door criterion [17, 18, 4, 5] based on the casual sufficiency assumption, and it is equivalent to the ignorability actually. (iv) It is difficult for POM to analyze indirect causation (mediation), thus POM typically relies on Baron-Kenny model [19, 20] for assistance. By contrast, it is easier for SCM by using front-door criterion [17, 18, 4, 5], that decomposes a causal pathway into two parts, i.e., from source variable to mediator and from mediator to target variable, and then conduct back-door adjustment twice if no other mediator.
Moreover, whether it is POM or SCM currently, they are both unable to model the spatio-temporal interactions between system units. E.g., the SUTVA requires no interaction between system units in POM. Moreover, SCM based on the traditional Bayesian network has no modeling for multivariate time series, thus the nonstationarity in time series is often neglected. Thus, this motivates us to find a common point to unify the two frameworks, and strengthen the inference for the spatio-temporal counterfactuals. Fortunately, our recent work [21, 22] has proposed a Spatio-Temporal Bayesian Networks (STBNs) model based on temporal assumption, that assumes the cause precedes or parallels the effect in time. And it is found that, the famous complex networks [23, 24, 25], a type of network dynamics to model spatio-temporal interactions, are a special case of STBNs.
Thus, we further conduct this work to build a spatio-temporal graphical counterfactual framework based on STBNs. The framework requires two components: (i) an STBNs used as the foundation of SCM, (ii) a Forward Counterfactual Inference Algorithm designed on the decomposability of the front-door criterion. Then we discuss how to get the two points, following the storyline as shown in Fig. 1. Thus, the main contributions of this work are as follows:
  1. 1.
    A survey is investigated to discuss the frameworks from POM to SCM, including their theoretical foundations and application approaches. This part is organized from Section 2 to Section 4.
  2. 2.
    A Forward Counterfactual Inference algorithm is designed here to recursively and autonomously infer counterfactuals through causal graphical languages. This part is presented as Algorithm 1 in Section 4.3.
  3. 3.
    An overview of the spatio-temporal graphical counterfactual framework is proposed to discuss Spatio-Temporal Bayesian Networks, its nonstationarity, and the relationship with complex networks. This part is organized in Section 5.
Figure 1: Towards spatio-temporal graphical counterfactuals.
2 Potential Outcome Model
The counterfactuals are defined as the unobserved potential outcomes in POM, as shown in Table 1. Here, $Y$ is the potential outcomes of the target variable. $T$ is treatment variable. $T=1$ represents being treated and $T=0$ represents being controlled, with a group of randomized control trials on $N$ units. $X$ represents covariates that have relationship with $Y$ and $T$, and are controlled for randomized trials. POM models the general process from causal estimands ($T$, $Y$ and $X$) to “Science” [26]. In this case, for different treatments $T$, only one potential outcome, $Y(1)$ or $Y(0)$, can be observed, and another one is unobserved, namely the counterfactual.
Table 1: Data for POM example.
   Units    Covariates    Potential outcomes    Treatment
   $i$    $X$    $Y(1)$    $Y(0)$    $T$
   1    $x_{1}$    $y_{1}(1)$    ??    11
   2    $x_{2}$    ??    $y_{2}(0)$    0
   3    $x_{3}$    ??    $y_{3}(0)$    0
   4    $x_{4}$    $y_{4}(1)$    ??    11
   5    $x_{5}$    $y_{5}(1)$    ??    11
   $\vdots$    $\vdots$    $\vdots$    $\vdots$    $\vdots$
   NN    $x_{N}$    ??    $y_{N}(0)$    0
2.1 Inferring Counterfactuals through POM
To infer the counterfactuals through POM, one common approach is matching, e.g., exact matching. For a unit $i$, one can search the samples with exactly matching covariates $X$ but treatment $T$ is different, and then estimate the counterfactual outcome $Y_{i}(1-T_{i})$. But it is difficult to conduct if the dimension of $X$ is high, that requires large number of samples to support the matching. Thus, approximate matching is commonly better, e.g., caliper matching [27, 28, 29, 30] and propensity score matching [11, 31, 32, 33]. They approximately search a group of similar samples for the target unit with one or multiple measurements, and finally calculate the counterfactuals.
Another common approach is data-driven imputation, that views the counterfactuals as missing values, and interpolates them by fitting the observational data. To realize it, linear or nonlinear regressions [34, 35, 36, 37] are commonly used to interpolate the missing values from the trained regressive models on the observational data. Different from this, tensor decomposition approaches [38, 39, 40] recover the missing values by decomposing and reconstructing the sparse data tensor. Moreover, deep generative models [41, 42, 43, 44, 45, 46, 47] are also used to model the uncertainty in data, and generate the missing values by randomly sampling. Note that, the generated missing values are not unique, but a group of values conforming to a random distribution. This is different compared to the regressive approaches and tensor decomposition approaches.
2.2 Foundational Assumptions of POM
Before using the above approaches on the framework of POM, four assumptions are necessary to be accepted. The first is SUTVA, which is defined as {assumption}[SUTVA [10]] In Table 1,
  1. 1.
    there is no interference between units, that is, neither $Y_{i}(1)$ nor $Y_{i}(0)$ is affected by what action any other unit received.
  2. 2.
    there is no hidden versions of treatments, that is, no matter how unit $i$ received treatment $T=1$, the outcome that would be observed would be $Y_{i}(1)$, and similarly for treatment $T=0$.
SUTVA guarantees the identical independence between the experimental units, and for unit $i$, the outcome $Y_{i}$ is only up to its treatment $T_{i}$, not the others’. However, this assumption is not always satisfied, e.g., in the case of social network, in which there are frequent interactions between units with varying time.
The second assumption is consistency, defined as {assumption}[Consistency [8, 9]] In Table 1, if the $i$-th unit is selected for a treatment $T_{i}$, the observed value of $Y_{i}$, neither $Y_{i}(1)$ nor $Y_{i}(0)$, is the same for all assignments of treatments to the other experimental units. Another form to describe the consistency assumption through expectation $\mathbb{E}[\cdot]$ is
$\mathbb{E}[Y(1)|T=1]=\sum_{i=1}^{N}y_{i}\times t_{i}=\mathbb{E}[Y|T=1],$ (1)
$\mathbb{E}[Y(0)|T=0]=\sum_{i=1}^{N}y_{i}\times(1-t_{i})=\mathbb{E}[Y|T=0],$
where $y_{1},\dots,y_{N}$ are the samples for $N$ units, and $t_{1},\dots,t_{N}\in\{0,1\}$ are the values of treatment variable $T$, as shown in Table 1. This is the same to the case of conditional expectation with $X$. Thus, the consistency assumption guarantees the missing counterfactual values in Section 2.1 can be interpolated by the other observational values actually.
The third assumption is positivity, defined as {assumption}[Positivity [11]] In Table 1, $0<P(T=1|X=x)<1$. This means for $X=x$, there are always some treatments randomly assigned with $T=0$ and $T=1$. Otherwise, the counterfactuals for $T=1$ could not be obtained if all units with $X=x$ are assigned with $T=0$, and this is the same to the counterfactuals for $T=0$.
The forth assumption is ignorability, defined as {assumption}[Ignorability [8, 9]] In Table 1, $T{\perp\!\!\!\perp}Y(1),Y(0)|X$ for each unit. Thus, the ignorability assumption implies that all the confounders have been observed in $X$, that is, no hidden confounder would interfere the observations of $Y(1),Y(0)$. Thus, the ignorability assumption is also called as unconfoundedness. This assumption are usually made to avoid the preference on experimental manipulation, that would guarantee the randomization in the experiments [48].
3 Structural Causal Model
Different from POM, SCM provides a type of graphical language to discuss causality intuitively, which avoids complex human-guided technical operations. An SCM is built on a set of variables, including endogenous variables $\mathbf{V}=\{X_{1},\dots,X_{n}\}$ and exogenous variables $\mathbf{U}=\{U_{1},\dots,U_{n}\}$. Note that, endogenous variables are observable from data, but exogenous variables are not. Moreover, a set of functions $\mathbf{F}=\{f_{1},\dots,f_{n}\}$ are also set to describe the functional relationship between these variables.
For example, as shown in Fig. 2, the SCM can be functionally described as
$f_{1}:$ $X_{1}=U_{1},$ (2)
$f_{2}:$ $X_{2}=aX_{1}+U_{2},$
$f_{3}:$ $X_{3}=bX_{1}+cX_{2}+U_{3},$
where $a=0.5,b=0.7,c=0.4$. $U_{1},U_{2},U_{3}$ are all additive noise, and they are mutually independent.
Note that, in an SCM, the functional relationship must conform to a directed-acyclic graphical constraint. The graph is early originated from the concept of Bayesian network [12, 13, 14, 15, 16]. And latter, it is also called as causal graphical model to highlight the causality [49, 17, 18, 4, 5]. And recently, it is also called as causal network in the field of causal discovery [50, 51, 52, 53, 54]. But actually, they are the same in the framework of SCM, and they are all directed acyclic graph (DAG) that represents Markovian knowledge. Thus, to avoid misunderstanding, we here use the name of Bayesian network uniformly, due to the fact that causal graphical model and causal network are not always a Bayesian network in Fig. 2. For example, Direct Cyclic Graph [55, 56, 57, 58], Markov network [59, 60], and Full Time Graph [6], they are also causal graph or causal network.
Figure 2: An example of SCM [5].
3.1 Pearl’s Causal Ladder [4, 5, 7]
To infer the counterfactuals through SCM, there are three steps:
  1. 1.
    Abduction: Infer the values of exogenous variables $\mathbf{U}$ from the observational data;
  2. 2.
    Action: Perform an intervention on SCM, e.g., $do(X_{2}=2)$ on SCM in Fig. 2, and then, $X_{2}$ would be assigned with value 22, and the arrows to $X_{2}$ would be modified, as shown in Fig. 3;
  3. 3.
    Prediction: Use the modified SCM to recalculate the counterfactuals of the target variable.
Figure 3: A diagram of Pearl’s causal ladder [5].
As shown in Fig. 3, we query “observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”. Obviously, this is a counterfactual question. Thus, follow the Pearl’s causal ladder, as shown in Fig. 3, we first infer the values of exogenous variables $\mathbf{U}=\{U_{1},U_{2},U_{3}\}$ from the observational data $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, as follows:
[…]
Then, $do(X_{2}=2)$ to obtain a modified SCM. And finally, recalculate the counterfactual of $X_{3}$, that is,
[…]
where $X_{3}(do(X_{2}=2))$ is different compared to the observational $X_{3}$.
The intervention, also named as $do$-operator [4], can be generalized into the form of probability, if modularity assumption is introduced as follow: {assumption}[Modularity [4]] If a set of variables $\{X_{j_{1}},\dots,X_{j_{p}}\}\subset\mathbf{V}$ is intervened, then for each variable $X\in\mathbf{V}$, it is obtained that
  1. 1.
    if $X\notin\{X_{j_{1}},\dots,X_{j_{p}}\}$, then $P(X|\mathbf{Pa}(X))$ remains unchanged. Here, $\mathbf{Pa}(X)$ is the predecessors of $X$ in Bayesian network, that is also called as causal parents.
  2. 2.
    if $X\in\{X_{j_{1}},\dots,X_{j_{p}}\}$, then $P(X=x|\mathbf{Pa}(X))=1$ if $x$ is the value set by intervention to $X$, otherwise, $P(X=x|\mathbf{Pa}(X))=0$.
This means, if a group of interventions are conducted on $X_{j_{1}},\dots,X_{j_{p}}$, the values of them would be fixed. The connections between them and their causal parents would be broken, and their causal parents would not affect them anymore in the modified SCM, just like in Fig. 3. Moreover, this also means that the probability distributions of the other variables that are not be intervened would not change.
3.2 Foundational Assumptions of Bayesian Network
As presented above, to clearly define the causality, every SCM is associate to a Bayesian network with directed-acyclic graphical constraint. Thus, the causality in Bayesian network determines the counterfactuals foundationally. Thus, we discuss the foundational assumptions of Bayesian network in the following contents.
First of all, a concept of d-separation can be defined on a Bayesian network $\mathcal{G}$ as
Definition 3.1 (d-separation [12, 13, 14, 15, 16]).
Let $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$ be three disjoint subsets of endogenous variables $\mathbf{V}=\{X_{1},\dots,X_{n}\}$, and let $\mathbf{p}$ be any path from a node in $\mathbf{X}$ to a node in $\mathbf{Y}$ regardless of direction. $\mathbf{Z}$ is said to block $\mathbf{p}$ if and only if there is a node $v\in\mathbf{p}$ satisfying one of the following items:
  1. 1.
    $v$ has v-structure (two nodes $a,b\in\mathbf{p}$ pointing to $v$, namely $a\rightarrow v\leftarrow b$), and neither $v$ nor its any descendants are in $\mathbf{Z}$;
  2. 2.
    $v$ in $\mathbf{Z}$ and $v$ does not have v-structure.
Then, $\mathbf{Z}$ d-separate $\mathbf{X}$ and $\mathbf{Y}$, denoted as $\mathbf{X}{\perp\!\!\!\perp}_{\mathcal{G}}\mathbf{Y}|\mathbf{Z}$, if $\mathbf{Z}$ blocks any $\mathbf{p}$.
Then, two assumptions, causal Markov (or Markov property) and faithfulness, are introduced as follows: {assumption}[Causal Markov [12, 13, 14, 15, 16]] Probability distribution $P$ is Markovian to a Bayesian network $\mathcal{G}$ if
$\mathbf{X}{\perp\!\!\!\perp}_{\mathcal{G}}\mathbf{Y}|\mathbf{Z}\Rightarrow\mathbf{X}{\perp\!\!\!\perp}\mathbf{Y}|\mathbf{Z},$ (5)
where $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$ are three disjoint subsets. {assumption}[Faithfulness [12, 13, 14, 15, 16]] Probability distribution $P$ is faithful to a Bayesian network $\mathcal{G}$ if
$\mathbf{X}{\perp\!\!\!\perp}\mathbf{Y}|\mathbf{Z}\Rightarrow\mathbf{X}{\perp\!\!\!\perp}_{\mathcal{G}}\mathbf{Y}|\mathbf{Z}$ (6)
for all disjoint subsets $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$. It is intuitive to understand the two assumptions, that is, the Bayesian network $\mathcal{G}$ has one-to-one correspondence to the independence of probability distribution $P$ in observational data. Thus, if they are satisfied, the joint distribution can be factorized according to the Bayesian network, as follow:
$P(X_{1},\dots,X_{n})=\prod_{i=1}^{n}P(X_{i}|\mathbf{Pa}(X_{i})),$ (7)
where $\mathbf{Pa}(X_{i})$ is the predecessors of $X_{i}$ in $\mathcal{G}$, that is also called as causal parents.
Moreover, a Bayesian network is also assumed to be causally sufficient, that is {assumption}[Causal Sufficiency [12, 13, 14, 15, 16]] Variables $\mathbf{V}$ are said to satisfy causal sufficiency if there is no hidden variable that is a common cause of two or more variables in $\mathbf{V}$. This means, if a Bayesian network (or an SCM) is causally sufficient, there is no hidden path connecting two endogenous variables in $\mathbf{V}$, and all variable information is collected sufficiently to support the network. Actually, the ignorability assumption in POM (see Assumption 2.2) also implies the causal sufficiency, and this assumption is weaker than ignorability. But note that, the causal sufficiency is hard to satisfy practically, because many real-world systems (e.g., economic system and climate system) are complex, thus, we could not collect all the information to describe the systems usually. Thus, some research works focus on this issue. For example, FCI (or Fast Causal Inference) algorithm and its variants [61, 62, 63, 64, 65, 66] are proposed to discover causality in the presence of hidden variables. And some other related works [67, 68, 69] investigate the causal discovery with soft intervention, that intervenes the SCM but does not change the network structure. There are also some related works [70, 71, 72, 73] proposed approaches to search the optimal and efficient adjustment set, that all its variables are observable, minimal and valid, and the removal of any of its variables would destroys the validity.
4 Differences between POM and SCM
In this section, the fundamental differences between POM and SCM are discussed, including (i) counterfactual falsifiability: POM assumes the positivity, thus the counterfactuals in POM can be falsified on observations, but SCM does not. The counterfactuals in SCM can be outside the observations, thus it is unfalsifiable. (ii) using back-door criterion to satisfy ignorability: POM requires ignorability to guarantee the unconfoundedness and correctness of counterfactual inference, but it is a complex human-guided operation. By contrast, SCM realizes unconfoundedness by the causal sufficiency and back-door adjustment, that are equivalent to the ignorability assumption. SCM facilitates counterfactual inference through graphical language. (iii) direct causation and indirect causation: POM struggles to process the indirect causation, and it typically relies on Baron-Kenny model. But in contrast, SCM has advantages to adjust the front-door path (that has indirect causation) through equivalently applying the back-door criterion twice. We further use this advantage to design the forward counterfactual inference algorithm to infer the spatio-temporal counterfactuals autonomously. The following are the details.
4.1 Counterfactual Falsifiability
As shown in Fig. 3, if we query to a POM, “Observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”, we must ensure there are at least two observational samples, and they have the same values $X_{1}=0.5$ but different values in $X_{2}$. That is, one is $X_{2}=1$, and another one is $X_{2}=2$, to satisfy the positivity assumption (see Assumption 2.2). And then, due to the consistency assumption (see Assumption 2.2), we can infer the counterfactuals through matching the two samples exactly. But this is different in SCM, we can infer the counterfactuals through the Pearl’s causal ladder, even if we only have a single sample, as shown in Section 3.1. Thus, how to falsify the counterfactuals?
It is a difficult question, because the counterfactuals are unobserved according to the definition (see “?” in Table 1), and we need at least two samples for comparison. In general, we cannot back to the past to falsify it with a different treatment, and we cannot reproduce the same experiments completely at different times. Thus, the inferred counterfactuals in SCM cannot be falsified without the assumptions of positivity and consistency, and this view is rejected ambiguously in Pearl’s book [7].
Another explanation is that, without the positivity, but with the consistency only in POM, the counterfactuals can be extrapolated by regressive approaches, and they are falsifiable. For example, as shown in Fig. 5, if there are $T$ and $Y$ only, and the observational data $(T,Y)=\{(0,0.5),(1,1),(2,2.5),(3,2)\}$ is collected, the linear regressive model can be fitted as $Y=0.5T+0.5$. Then, we query “if had done $T=4$, what would Y=?Y=?”, the counterfactual answer is $Y=2.5$, and the answer can be falsified by the regressive function because it does not beyond the observation. Thus, the consistency assumption actually claims that the counterfactuals can be falsified by the observations. This is also true for SCM, because SCM also uses the regressive approaches to infer counterfactuals.
Moreover, there may be some confusions regarding the preset Bayesian network (see Fig. 2). To build an SCM, a Bayesian network is necessary to be known first. However, how to discover a probabilistic network from a single sample? This is usually impossible. A reasonable explanation is that, the network structure can be discovered from the system mechanisms (e.g., physical mechanisms, communication connections), or graphical knowledge learned from other domains, instead of probability. However, if it is accepted, the definition of causality could be broader and more ambiguous. This is why probabilistic and causal Bayesian network are often indistinguishable, as stated in Pearl’s book [7].
Thus, actually, SCM provides a way to think the counterfactuals, and it emphasizes the thinking way is counterfactual, but not the thinking result, because the result cannot always be falsified without the positivity and consistency assumptions. And on the contrary, POM emphasizes the thinking result is counterfactual. This is a fundamental difference in the problem of counterfactual falsifiability.
4.2 Ignorability and Back-door Criterion
In the view of statistics, the ignorability assumption (see Assumption 2.2) is equivalent in calculation to the causal sufficiency assumption (see Assumption 3.1) with a back-door criterion. The accurate definition of back-door criterion can be found in these papers and books [17, 18, 4, 5]. Here we present another from with respect to the Table 1, defined as
Definition 4.1 (Back-door Criterion [17, 18, 4, 5]).
In Table 1, the covariate $X$ is said to satisfy back-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfy
  1. 1.
    no variable in $X$ is a descendant of $T$;
  2. 2.
    $X$ blocks every path between $T$ and $Y$ that contains an arrow into $T$.
Figure 4: A diagram of data extrapolation.
Figure 5: A diagram of back-door path.
Thus, for $X,Y,T$, only one structure can satisfy the back-door criterion, as shown in Fig. 5. And in the back-door path, it is obtained that
$P(Y|do(T=1))$ (8)
$=$ $\sum_{x}P(Y|do(T=1),X=x)P(X=x|do(T=1))$
$=$ $\sum_{x}P(Y|T=1,X=x)P(X=x),$
which is the same in case of $do(T=0)$ [17, 18, 4, 5]. Then, if the causal sufficiency is satisfied, that is, no hidden variable can interfere $X,Y,T$, the expectation of potential outcomes in Table 1 can be calculated by
$\mathbb{E}[Y|do(T=1)]$ (9)
$=$ $\sum_{y}y\times P(Y=y|do(T=1))$
$=$ $\sum_{y}\sum_{x}y\times P(Y=y|T=1,X=x)P(X=x)$
$=$ $\sum_{x}P(X=x)\sum_{y}y\times P(Y=y|T=1,X=x)$
$=$ $\mathbb{E}[\mathbb{E}[Y|T=1,X]].$
This is the same in POM, as follow:
$\mathbb{E}[Y(1)]$ (10)
$=$ $\mathbb{E}[\mathbb{E}[Y(1)|X]]$ $/*Law~of~full~expectation*/$
$=$ $\mathbb{E}[\mathbb{E}[Y(1)|T=1,X]]$ $/*Ignorability*/$
$=$ $\mathbb{E}[\mathbb{E}[Y|T=1,X]],$ $/*Consistency*/$
which is the same in case of $do(T=0)$. Thus, with comparing Eq. (9) and Eq. (10), $\mathbb{E}[Y|do(T=1)]=\mathbb{E}[Y(1)]$, and $\mathbb{E}[Y|do(T=0)]=\mathbb{E}[Y(0)]$, in the two frameworks, but they start from different assumptions. Thus, the ignorability assumption can be decomposed as the causal sufficiency assumption and the back-door criterion, and the causal sufficiency is relatively trivial to satisfy. Actually, the ignorability is more likely to be a technical assumption, that needs many controlled trials to deconfound, while the back-door path is clear to search in a Bayesian network. A potential and more difficult technical issue is to discover the network structure accurately.
4.3 Direct and Indirect Causation
Suppose that there is a causal ordered pathway from one treatment variable to another outcome variable. Then, if there are one or more variables between the two endpoints, the pathway is called as indirect causation, otherwise, it is called as directed causation. Actually, the indirect causation is also called as mediating effect, that has been discussed deeply in Rubin’s papers [74, 75, 76]. Here, Baron-Kenny model [19, 20] is introduced to present mediating analysis for potential outcome with graph. As shown in Fig. 7, $T\rightarrow Y$ is direct causation, and $T\rightarrow X\rightarrow Y$ is indirect causation.
Figure 6: A diagram of mediating analysis [19]. Here, $a,b,c$ are coefficients and $U_{X},U_{Y}$ are biases. $b^{\prime}$ and $U_{total}$ are the coefficient and bias respectively for total effect.
Figure 7: A diagram of front-door adjustment [5]. Here, $T$ is a treatment variable, and $Y$ is a target variable. $X_{1}$ and $X_{2}$ are both covariates.
To analyse indirect causation, causal-steps approach [19] is proposed, through building three linear regressive models, as shown in Fig. 7. Suppose that the regressive models are fitted sufficiently, the indirect causation would be detected if satisfy
  1. 1.
    coefficients $a,b$ and $c$ are significant;
  2. 2.
    $|b|<|b^{\prime}|$.
Another approaches are to test the significance of $H_{0}:b^{\prime}-b=0$ [77, 78, 79]. Or, to test the significance of $H_{0}:ab=0$ [80, 81]. If the correct regressive models are built, the counterfactuals can be calculated, just like SCM in Fig. 2. Moreover, they are all regression-based approaches, that restrict the usage in large multivariate datasets, because one needs to test whether a covariate is a confounder or a mediator, and the test process would be time-consuming seriously.
But on the contrary, the test process can be faster with Bayesian network in an SCM. Here, we propose an algorithm to realize forward counterfactual inference as shown in Algorithm 1. To clearly demonstrate the principle of Algorithm 1, we first introduce the front-door criterion as follows:
Definition 4.2 (Front-door Criterion [17, 18, 4, 5]).
The covariate $X$ is said to satisfy front-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfy
  1. 1.
    $X$ intercepts all directed paths from $T$ to $Y$;
  2. 2.
    there is no back-door path from $T$ to $X$;
  3. 3.
    all back-door paths from $X$ to $Y$ are blocked by $T$.
Then, a classic example is provided to show the front-door adjustment through the theorem 3.4.1 in book [5], as also shown in Fig. 7 here. If given a treatment variable $T$ and a target variable $Y$, with two covariates $X_{1}$ and $X_{2}$. Then one can adjust the distribution $P(Y|do(T))$ with the front-door criterion as follows:
$P(Y|do(T))$ $=\sum_{x_{1}}P(X_{1}|do(T))P(Y|do(X_{1}))$ (11)
$=\sum_{x_{1}}P(X_{1}|T)\sum_{t^{\prime}}P(Y|X_{1},T^{\prime})P(T^{\prime})$
Thus, we can conclude from this equation:
  1. 1.
    If there are one or multiple ordered causal pathways from $T$ to $Y$ and there is at least one mediator on the ways, then the causal pathway can be decomposed into two causal pathways through that mediator, e.g., $T\rightarrow X_{1}$ and $X_{1}\rightarrow Y$ in Eq. (11).
  2. 2.
    If there is no mediator on all causal pathways, then adjustment using back-door criterion. Of course, if there is no back-door path, no need to adjust.
Thus, the front-door principle is non-atomic and decomposable. And with this more foundational principle, we propose the forward counterfactual algorithm to autonomously apply this adjustment principle, as presented in Algorithm 1. The algorithm can be conducted on arbitrary variable set $\{X_{1},\dots,X_{n}\}$ that satisfied causal sufficiency. And we also provide an example to show the inference process as shown in Fig. 8.
In Algorithm 1, the approaches to realize the codes in lines 1 and 3 can be various, e.g., some variants of depth-first-search algorithm and breadth-first-search algorithm [82]. Thus, their time complexities are both $\mathcal{O}(|\mathbf{V}|+|\mathbf{E}|)$, if a Bayesian network $\mathcal{G}=(\mathbf{V},\mathbf{E})$ is given, where $\mathbf{V}$ is the node set and $\mathbf{E}$ is the edge set. Meanwhile, if the computational cost of factorization and adjustment are overlooked, the time complexity of the code in line 9 in Algorithm 1 is $\mathcal{O}(|\mathbf{V}|)$, because we only need to traverse all mediators in each pathway forward, and this process can be performed in parallel. Thus, the total time complexity of Algorithm 1 is $\mathcal{O}(|\mathbf{V}|^{2}+|\mathbf{V}||\mathbf{E}|)$. Thus, if the network structure of the Bayesian network is large, and it can be discovered accurately, the efficiency of SCM to infer counterfactuals would be better with polynomial time complexity.
5 Spatio-Temporal Graphical Counterfactuals
Then for the examples of counterfactual questions at the beginning of this article, that is, if a different investment strategy had been implemented, would we obtain higher returns? And, in computer networks, what changes would occur in network load if a node’s configuration had never been changed? And, would someone still purchase the product even if they had never been shown the advertisement? In these real-world scenarios, interactive behaviors with lagged causal effects are widely-existing. POM cannot answer these questions, because POM does not allow the interactive behaviors between experimental units (see SUTVA in Assumption 2.2), even with some temporal quasi-experimental approaches (e.g., Regression Discontinuity Design [83, 84] and Differences in Differences [85, 86]). And this is the same to SCM, because the foundational Bayesian network is a DAG that also does not allow mutual connections between network nodes. Thus, to answer these counterfactual queries for interactive units and lagged causal effects, a concept of spatio-temporal graphical counterfactuals is proposed here.
Input: A Bayesian network $\mathcal{G}$ built on $\mathbf{V}=\{X_{1},\dots,X_{n}\}$ with one target $X_{i}$ and multiple sources $X_{j_{1}},\dots,X_{j_{p}}$.
Output: Distribution $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$.
1
2Search one or multiple ordered causal pathways in $\mathcal{G}$ from the sources $X_{j_{1}},\dots,X_{j_{p}}$ to the target $X_{i}$;
3if $X_{j_{1}},\dots,X_{j_{p}}$ are all the direct causation for $X_{i}$ then
4   Search back-door paths (common causes) starting from $X_{j_{1}},\dots,X_{j_{p}}$ and $X_{i}$ reversely in $\mathcal{G}$.
5 if no back-door path within them then
6    return $P(X_{i}|X_{j_{1}},\dots,X_{j_{p}})$;
7 else
8      For every back-door path, select a group of direct predecessors, $\mathbf{W}\subset\mathbf{V}$, and then, according to back-door criterion (see Definition 4.1),
9    return $\sum_{w}P(X_{i}|X_{j_{1}},\dots,X_{j_{p}},W)P(W)$;
10
11
12Otherwise, for $q$ pathways starting from $X_{j_{p}}$, select $q$ first mediators $X_{j_{p_{1}}},\dots,X_{j_{p_{q}}}$. Then, a group of source-mediator pairs, e.g., $(X_{j_{p}},X_{j_{p_{q}}})$, can be obtained;
13Factorize $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$ as
$P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$ (12)
$=$ $\sum_{x_{j_{1_{1}}:j_{p_{q}}}}P(X_{i}|do(X_{j_{1_{1}}}),\dots,do(X_{j_{2_{1}}}),\dots,do(X_{j_{p_{q}}}))$
$\times\prod_{p}\prod_{q}P(X_{j_{p_{q}}}|do(X),X\in\mathbf{Pa}(X_{j_{p_{q}}})\subseteq\{X_{j_{1}:j_{p}}\}),$
where $\mathbf{Pa}(X_{j_{p_{q}}})$ is the causal parents of $X_{j_{p_{q}}}$;
14Solve each distribution $P(\cdot|do(\cdot),\dots,do(\cdot))$ in Eq. (12) by Algorithm 1 recursively.
return $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$.
Algorithm 1 Forward Counterfactual Inference
Figure 8: An example of Forward Counterfactual Inference algorithm. Here, $X_{i}$ is the same as $X_{i}^{\prime}$ to distinguish two recursions.
Table 2: Comparisons of various causal graphs.
Temporal Bayesian Network Dynamical Bayesian Networks Full Time Graph Spatio-Temporal Bayesian Networks
High-order causation $\times$ $\times$ $\checkmark$ $\checkmark$
Instantaneous causation $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Nonstationary causation $\times$ $\checkmark$ $\times$ $\checkmark$
5.1 Spatio-Temporal Bayesian Networks
Before the buildings of SCM, actually, various causal graphical models have been proposed to model the spatio-temporal causality. As shown in Fig. 9, Temporal Bayesian Network [59, 87] are proposed to model the momentary causal dependency in a first-order Markov process. Further, Dynamical Bayesian Networks [59] are proposed to represent a combination of multiple Temporal Bayesian Networks. Full Time Graph [6] is proposed to summarize the temporal Bayesian dependency at full timestamps. Note that, Full Time Graph allows the instantaneous causal effects between two units (or variables) in a high-order Markov process, but it only allows strictly stationary causality, that is, the causal dependencies do not change with time. Moreover, Full Time Graph is proved to uniquely decomposed into multiple high-order Temporal Bayesian Networks, named UCN, if the instantaneous causal effects is not allowed [21]. The comparisons of these causal graphs are shown in Table 2.
Figure 9: Diagrams of various causal graphs.
Thus, based on the characteristics of these causal graphs, STBNs are defined as a group of DAGs with different timestamps, as shown in Fig. 9. In STBNs, intuitively, the nodes are not equivalent to the variables, but represent the state of the variables at some time steps. With the assumptions of causal Markov (see Assumption 5) and faithfulness (see Assumption 6), the joint distribution of all nodes can be factorized as
$P(X_{1,t},\dots,X_{n,t},X_{1,t-1},\dots,X_{n,t-T+1})$ (13)
$=$ $\prod_{\tau=0}^{T-1}\prod_{i=1}^{n}P(X_{i,t-\tau}|\mathbf{Pa}(X_{i,t-\tau})),$
where $X_{1},\dots,X_{n}$ are the variables, and the timestamps are in range of $0\sim T-1$. Theoretically, infinite time steps are allowed in STBNs, but it is usually finite in practice.
To guarantee the global and unique DAG in STBNs, a temporal assumption is introduced as {assumption}[Temporal Assumption] In STBNs, the cause node precedes or parallels the effect node in time. This assumption does not allow the directed cause-effect pairs from now to the past, e.g., $X_{i,t-\tau_{1}}\rightarrow X_{j,t-\tau_{2}},\tau_{1}<\tau_{2}$. And this guarantees the functionally identifiable network structure without Markov equivalence class [6, 21, 88]. Moreover, the causal sufficiency assumption (see Assumption 3.1) is also needed to guarantee the unconfoundedness.
Then, if STBNs are given, the spatial-temporal counterfactuals can be inferred through Algorithm 1, because the global and local STBNs are both DAG. Compared with the naive Bayesian networks, STBNs have the spatio-temporal concept, and the inferring process is naturally forward along with time. This enables to define the momentary interventions for every time steps, and all the interventions would work sequentially, instead of at the same time, like in Algorithm 1. If fully consider this property, the inference algorithm can be faster.
5.2 Nonstationarity of STBNs
Nonstationarity is a significant topic in researches on time series. Actually, however, the reasons inducing it are likely to be different from the view of causality. Three of the most currently popular reasons are:
  1. 1.
    Trending: The time series have ascending or descending trending for long terms.
  2. 2.
    Time-varying Variance: The variances of additive noise are time-varying, that is also called as heterogeneous causality in papers [89, 90].
  3. 3.
    Nonstationary Causality: The structures of STBNs change over time, and at least once, as shown in Fig. 10.
Figure 10: Diagrams of nonstationary causality and heterogeneous causality. Here, $C_{t-\tau}=t-\tau,\tau=0,\dots,T-1$ are usually assumed, thus, they are observable, instead of hidden variable.
For the trending time series, an effective approach is to calculate differences of time series in preprocessing. This approach enables to obtain the differential time series, that is the observations of system dynamics, like Eq. (14). This implied, the dynamics is stable during the period of time, and the structures of STBNs do not change, or in another word, the differential time series are stationary. Thus, if the dynamics equations, e.g., Eq. (14), are not equal to zero, the causal links in STBNs are identifiable through the Jacobian matrix of dynamics [91, 92, 93, 94].
For the case of time-varying variance, the causal identifiability is up to the mutually independent exogenous additive noises [88]. That is, if there is no causal connection between two additive noise nodes, the causation is identifiable. Thus, an effective approach is to relate the time-varying noises with a common timestamp, and the timestamps at all time steps are observable confounders $C_{t-\tau},\tau=0,\dots,T-1$ [95, 89, 96, 90]. As shown in Fig. 10, there are three causal pathways, that is, $C_{t-1}\rightarrow U_{4,t-1}\rightarrow X_{4,t-1}$, $C_{t-1}\rightarrow C_{t}\rightarrow U_{3,t}\rightarrow X_{3,t}$, and $C_{t-1}\rightarrow C_{t}\rightarrow U_{4,t}\rightarrow X_{4,t}$. Thus, if $C_{t}$ and $C_{t-1}$ are hidden, the spurious connection between $X_{4,t-1}$ and $X_{3,t}$ would be detected, like the dotted line in Fig. 10. However, $C_{t}$ and $C_{t-1}$ are usually known as the timestamp $t$ and t1t-1, thus, $C_{t}$ and $C_{t-1}$ block the fork paths between $X_{4,t-1},X_{3,t}$ and $X_{4,t}$. Then, the noise variables $U_{4,t-1},U_{3,t}$ and $U_{4,t}$ are mutually independent. Thus, it is causally identifiable. Note that, however, the premise to do so is that any hidden confounders in the case of time-varying variance can be written as smooth or non-smooth functions of timestamps [90].
Nonstationary causality is challenging but critical, especially for the case of frequently changing structures. To identify the nonstationary causality, one approach is to use sliding windows to identify causation within different time periods, as shown in Fig. 10. For each sliding window, it is assumed that the causal structure is invariant, and then, a group of samples would be input to identification algorithms, e.g., [97, 98, 99, 100, 21, 22], to identify a invariant structure during this period. However, if the network changes are frequent, the samples used for identification during a short-term period would be few, but in contrast, we usually need sufficient samples to accept the significance of causal connections. Thus, another reasonable approach is to identify the change points first, and then, identify the causality with the corresponding periods [101, 102]. However, this still does not work in the extreme case of frequently changing, because the samples between two change points may be few. Moreover, there are some other approaches to view the nonstationary causality as a probabilistic normalized flow, and then use variational Bayesian inference to estimate it globally [103, 104, 105].
5.3 STBNs and Complex networks
Here, we suppose STBNs model has been built, and it is associated with interactive time-series process of multivariate (or multiple units). Then, we further introduce complex network, another type of network model, to view the causal interactions in complex system. The network dynamics for variables $X_{i},\dots,X_{n}$ can be defined as
$\frac{dX_{i,t}}{dt}=F(X_{i,t})+\sum_{j=1}^{n}A_{ij}G(X_{i,t},X_{j,t}),$ (14)
where $F(\cdot)$ is the self-governed function of $X_{i}$, and $G(\cdot,\cdot)$ is the interactive function for node pairs $X_{i}$ and $X_{j}$. Moreover, $A_{ij}\geq 0$ is the connection between $X_{i}$ and $X_{j}$. Thus, if Eq. (14) is discretized, we can find it the same as first-order STBNs on stationary time series, as shown in Fig. 11.
Figure 11: A diagram of discretized network dynamics. Here, $A$ is the adjacency matrix of the first-order STBNs, and $A_{ii}^{\prime},i=1,\dots,n$ represents the joined causal effect of $F$ and $G$.
Thus, the first-order STBNs have some good properties if network dynamics have. For example, if we want to know how to affect a node in STBNs by “intervening” another node, the practical inference steps may not be long, due to the small-world property [106, 107] or the scale-free property [108] of complex networks. Moreover, the counterfactual outcomes would not change because network systems have tolerance from external “interventions” [109, 110]. We would also know the synchronization [111, 112, 113, 114], controllability [115, 116, 117, 118, 119, 120, 121, 122], resilience [23, 123] of STBNs. We would also be able to sparsely identify the network structure even if the network size is large [24, 25], because the independence are equivalent to predictability in STBNs [21]. Also, the network size of STBNs can be reduced, because many complex networks can be simplified and still provide an insightful description of the causality of interest [124, 125]. This is meaningful to design counterfactual inference algorithm on smaller networks, and then obtain higher inference efficiency.
6 Conclusion
In this work, we mainly focus on the spatio-temporal graphical counterfactuals, and organize an overview for it to discuss its theoretical foundations and application approaches. To discuss the theoretical foundations, a survey is investigated, and the definition of counterfactuals is defined based on the concept of potential outcome in the framework of POM firstly. And then, the framework of SCM is also introduced to infer counterfactuals through graphical languages, which are equivalent to the counterfactuals in POM. Further, to infer counterfactuals on intelligent machine autonomously, a Forward Counterfactual Inference algorithm is designed in this work, and it is able to recursively solve the counterfactual probability distribution, $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$, with polynomial time complexity if multi-nodes interventions are conducted on $X_{j_{1}},\dots,X_{j_{p}}$. With this algorithm, the spatio-temporal graphical counterfactuals can be inferred if have two elements, that is, the network structure of STBNs and the corresponding algorithm to identify the network structure. This work not only discusses the identification algorithms in various nonstationary cases, but also discusses the feasibility to improve the algorithm efficiency from the view of complex networks.
[…]
This work is supported by the National Key R&D Program of China under Grant No. 2022ZD0120003, the National Natural Science Foundation of China under Grant Nos. 62233004, 62273090, 52525204, 52572354, the Jiangsu Provincial Scientific Research Center of Applied Mathematics under Grant No. BK20233002, the US National Science Foundation under Grant No. 2047488 and the Rensselaer-IBM AI Research Collaboration.